OpenStudy (anonymous):

Francisco has a savings account balance of $2,033.88. The interest rate on the account is 2.9% compounded monthly. If he opened the account nine years ago, what was the value of his initial deposit?

OpenStudy (kropot72):

\[2033.88=P(1+\frac{0.029}{12})^{(9\times 12)}\] \[P=\frac{2033.88}{(1.00241667)^{108}}\]